# On the generalized coloring numbers11footnotemark: 1

### v1

@@ 1 Introduction
Several generalizations of the coloring number have been studied in the literature, including the arrangeability by Chen and Schelp [23] (connected to Ramsey problems), the game chromatic number by Faigle et al. [63], the game coloring number by Zhu [157], the rank by Kierstead [91] (related to the game chromatic number), and the admissibility of a graph by Kierstead and Trotter [92] (used to bound both the arrangeability and the game chromatic number). These parameters are yet generalized by the following three series of parameters, called the generalized coloring numbers. For a positive integer $r$, the strong $r$-coloring number $\operatorname{col}_{r}(G)$ and weak $r$-coloring number $\operatorname{wcol}_{r}(G)$ were introduced by Kierstead and Yang [93], and the $r$-admissibility $\operatorname{adm}_{r}(G)$ was introduced by Dvořák [50]. All of these generalizations rely on linear orders and certain back-connectivity properties (formal definition will be given below).
Intuitively, the generalized coloring numbers capture structural properties not only of subgraphs of a graph, as degeneracy does, but also of its bounded-depth minors. Bounded-depth minors are the key concept in the theory of bounded expansion and nowhere dense graph classes, introduced by Nešetřil and Ossona de Mendez. The (non trivial) connection between the generalized coloring numbers and classes with bounded expansion was made by Zhu [158], and this connection naturally extends to nowhere dense classes as observed by Nešetřil and Ossona de Mendez [116].
@@
Bounded expansion and nowhere dense classes are intimately linked to first-order logic and many of the developments were driven by the question to solve the first-order model checking problem on these classes. Structural decompositions derived from the generalized coloring numbers lead to quantifier-elimination, and as a consequence to efficient model checking as shown by Dvořák, Král’ and Thomas [53], and enumeration and query counting as shown by Kazana and Segoufin [89], for first-order logic on classes with bounded expansion. Game characterizations and sparse neighborhood covers lead to tractable first-order model checking on nowhere dense classes as shown by Grohe et al. [77]. In fact, for monotone (that is, subgraph-closed) classes of graphs nowhere denseness turns out to be the dividing line between tractable and intractable first-order model checking. Hence, the classification of tractable first-order model checking on monotone graph classes is complete and leaves the question of a classification of tractability on hereditary graph classes.
@@
Figure 3: The map of the (structurally) sparse universe.
@@ 2.1 Treewidth and the strong r-coloring numbers
Treewidth, as a measure of the global connectivity of a graph, plays an important role in graph theory. Treewidth and tree decompositions were introduced by Robertson and Seymour [136] as part of their graph minors project, however, several strongly related or equivalent notions were invented independently [10, 79]. We refer to [13, 81] for surveys on various characterizations of treewidth and equivalent parameters and to [14, 28] for surveys on algorithmic applications.
@@
A tree decomposition of a graph $G$ is a pair $\mathcal{T}=(T,\mathrm{Bag})$, where $T$ is a tree and $\mathrm{Bag}\colon V(T)\to\mathcal{P}(V(G))$ is a mapping that assigns each node $x$ of $T$ to its bag $\mathrm{Bag}(x)\subseteq V(G)$ so that the following conditions are satisfied:
@@
In the following, it will be convenient to assume that $T$ is a rooted tree. A rooted tree is a tree $T$ with a designated root vertex $r$. Assigning a root imposes the standard ancestor/descendant relation $\leq_{T}$ in $T$ where the root is the $\leq_{T}$ minimal element: a node $v$ is a descendant of all the nodes that appear on the unique path leading from $v$ to the root. Note that we treat a vertex as an ancestor and descendant of itself. The ancestors and descendants of  $v$ excluding $v$ are its true ancestors and true descendants, respectively. The parent of a non-root node is its closest true ancestor. A rooted forest $F$ is a disjoint union of rooted trees. We write $\leq_{F}$ for the partial order that coincides with $\leq_{T}$ on each rooted tree in $F$.
@@
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a tree decomposition of $G$. This decomposition induces a quasi-order on $V(G)$ as follows. For $v\in V(G)$, let $x(u)$ be the unique $\leq_{T}$-minimal node of $T$ with $v\in\mathrm{Bag}(x)$. Define the quasi-order $\leq_{\mathcal{T}}$ on $V(G)$ by $u\leq_{\mathcal{T}}v$ if and only if $x(u)\leq_{T}x(v)$. Any linearization of $\leq_{\mathcal{T}}$ has very nice properties, which we will exploit later. In fact, we can associate a width to any linear order of $V(G)$ and thereby provide a characterization of treewidth.
@@
Let $G$ be a graph and let $\pi=(v_{1},\ldots,v_{n})$ be an order of $V(G)$. We construct the graph $G^{\pi}_{\textit{fill}}$, called the fill-in graph of $G$ with respect to the order $\pi$, by the following fill-in procedure. We let $G_{n}\coloneqq G$. Now, for $i=n-1,\ldots,1$, assume that the graph $G_{i+1}$ has been constructed. We define $G_{i}$ as the graph obtained from $G_{i+1}$ by adding all edges between non-adjacent neighbors of $v_{i+1}$ that are smaller than $v_{i}$. Finally, we let $G^{\pi}_{\textit{fill}}\coloneqq G_{1}$. It is immediate from the definition that $\pi$ is a perfect elimination order of $G^{\pi}_{\textit{fill}}$.
@@
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is strongly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}v$ for all internal vertices $w\in V(P)$. The set of vertices strongly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{SReach}_{r}[G,\pi,v]$. We define
@@ 2.2 Treedepth and the weak r-coloring numbers
Let $G$ be a graph. The treedepth $\operatorname{td}(G)$ of $G$ is the minimum depth of a rooted forest $F$ on the same vertex set as $G$ such that whenever $uv\in E(G)$, then $u\leq_{F}v$ or $v\leq_{F}u$.
@@
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is weakly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}u$ for all internal vertices $w\in V(P)$. The set of vertices weakly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{WReach}_{r}[G,\pi,v]$. We define
@@ 2.4 Logic and transductions
A variable $x$ not in the scope of a quantifier is a free variable (we do not consider formulas with free set second-order variables). A formula without free variables is a sentence. The quantifier rank $\mathrm{qr}(\varphi)$ of a formula $\varphi$ is the maximum nesting depth of quantifiers in $\varphi$. A formula without quantifiers is called quantifier-free.
@@
For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of colored graphs with bounded treedepth there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear-time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
@@ 3 General bounds for the generalized coloring numbers
In this section, we present lower and upper bounds for the generalized coloring numbers. It turns out that the numbers are intimately linked to the density of bounded-depth minors and topological minors. These are the central notions in the theory of bounded expansion and nowhere dense graph classes, which were introduced by Nešetřil and Ossona de Mendez [114, 116]. After introducing these concepts, we present simple lower bounds for the generalized coloring numbers in terms of the density of depth-$r$ minors, and depth-$r$ topological minors. Throughout this work when $r$ is not explictly quantified it denotes a non-negative integer.
@@ 3.1 Graph expansion and simple lower bounds
The following notions of bounded depth minors and bounded depth topological minors are the fundamental definitions in the theory of bounded expansion and nowhere dense graph classes introduced by Nešetřil and Ossona de Mendez [114, 116].
@@
Hence, expansion and topological expansion are functionally equivalent. The key observation for the proof is that in the branch set of a high-degree vertex in a bounded depth minor model, we will always find a vertex of large degree. We can use such a vertex as the branching vertex of a high-degree vertex in a topological minor model.
@@
Classes of bounded expansion are classes of sparse graphs by definition. It turns out that also nowhere dense classes are sparse.
@@ 3.2 Bounding r-admissibility
In view of Lemma 2.21, the algorithm of Theorem 3.15 also gives linear time computable approximations for the strong $r$-coloring numbers and weak $r$-coloring numbers. It was shown by Grohe et al. [76] that for $r\geq 3$, computing exact values for $\operatorname{wcol}_{r}(G)$ is NP-complete. Similarly, it is shown by Dvořák [48] that deciding whether a graph contains a depth-$1$ minor of density at least $d$ is NP-complete for $d\geq 2$ and by Muzi et al. [107] that deciding whether a graph contains a $1$-subdivision of density at least $d$ for $d\geq 2$ is also NP-complete.
Problem 3.16.
Let $r\geq 2$ be a positive integer. Does there exist an algorithm that, given an $n$-vertex graph $G$, computes $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ in time $n^{f(r,\operatorname{col}_{r}(G))}$, or $n^{f(r,\operatorname{wcol}_{r}(G))}$, for some function $f$?
@@ 3.4 Transitive fraternal augmentations
If $w$ is an $r$-fraternity function for every value of $r$, then $w$ is called a fraternity function.
@@ 4 Tree-decomposable graphs
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a rooted tree decomposition of $G$. For a non-root node $x$ with parent $y$, we define the adhesion of $x$ as $\mathrm{Adh}(x)=\mathrm{Bag}(x)\cap\mathrm{Bag}(y)$. If $x$ is the root, then we set $\mathrm{Adh}(x)=\emptyset$ by convention. The adhesion of the tree decomposition $\mathcal{T}=(T,\mathrm{Bag})$ is the maximum size of an adhesion set in $\mathcal{T}$, i.e., $\max_{x\in V(T)}|\mathrm{Adh}(x)|$. The torso of a node $x$ is the graph $\mathrm{Torso}(x)$ on vertex set $\mathrm{Bag}(x)$ where two vertices $u,v\in\mathrm{Bag}(x)$ are adjacent if and only if $uv\in E(G)$ or if there exists $y\neq x$ such that $u,v\in\mathrm{Bag}(y)$. Equivalently, $\mathrm{Torso}(x)$ is obtained from $G[\mathrm{Bag}(x)]$ by turning the adhesions of $x$ and of all children of $x$ into cliques. The margin of a node $x$ is the set $\mathrm{Margin}(x)=\mathrm{Bag}(x)\setminus\mathrm{Adh}(x)$. We say that $\mathcal{T}$ is a tree decomposition over a class $\mathscr{D}$ of graphs if $\mathrm{Torso}(x)\in\mathscr{D}$ for every node $x$ of $T$.
@@ 4.1 H-minor-free graphs
Let $G$ be a graph. A partition $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ of $V(G)$ is a connected partition if each $V_{i}$ induces a connected subgraph of $G$. We denote the minor of $G$ obtained from $G$ by contracting each $V_{i}$ to a single vertex $v_{i}$ by $G/\mathcal{P}$. We inherit all the above definitions for connected partitions by identifying a part $V_{i}$ with the vertex $v_{i}$ in (the ordered graph) $G/\mathcal{P}$.

### v3

@@ 1 Introduction
Several generalizations of the coloring number have been studied in the literature, including the arrangeability by Chen and Schelp [24] (connected to Ramsey problems), the game chromatic number by Faigle et al. [64], the game coloring number by Zhu [157], the rank by Kierstead [92] (related to the game chromatic number), and the admissibility of a graph by Kierstead and Trotter [93] (used to bound both the arrangeability and the game chromatic number). These parameters are yet generalized by the following three series of parameters, called the generalized coloring numbers. For a positive integer $r$, the strong $r$-coloring number $\operatorname{col}_{r}(G)$ and weak $r$-coloring number $\operatorname{wcol}_{r}(G)$ were introduced by Kierstead and Yang [94], and the $r$-admissibility $\operatorname{adm}_{r}(G)$ was introduced by Dvořák [53]. All of these generalizations rely on linear orders of the vertices and certain back-connectivity properties (formal definition will be given below).
Intuitively, the generalized coloring numbers capture structural properties not only of subgraphs of a graph, as degeneracy does, but also of its bounded-depth minors (see Figure 2). Bounded-depth minors are the key concept in the theory of bounded expansion and nowhere dense graph classes, introduced by Nešetřil and Ossona de Mendez. The (non trivial) connection between the generalized coloring numbers and classes with bounded expansion was made by Zhu [158], and this connection naturally extends to nowhere dense classes as observed by Nešetřil and Ossona de Mendez [114].
@@
Bounded expansion and nowhere dense classes are intimately linked to first-order logic and many of the developments were driven by the question to solve the first-order model checking problem on these classes. Structural decompositions derived from the generalized coloring numbers lead to quantifier-elimination, and as a consequence to efficient model checking as shown by Dvořák, Král’ and Thomas [55], and enumeration and query counting as shown by Kazana and Segoufin [90], for first-order logic on classes with bounded expansion. Game characterizations and sparse neighborhood covers lead to tractable first-order model checking on nowhere dense classes as shown by Grohe et al. [78]. In fact, for monotone (that is, subgraph-closed) classes of graphs nowhere denseness turns out to be the dividing line between tractable and intractable first-order model checking. Hence, the classification of tractable first-order model checking on monotone graph classes is complete and leaves the question of a classification of tractability on hereditary graph classes. Figure 3 shows an inclusion diagram of relevant graph classes and the current frontiers of tractable model checking.
@@
Figure 3: The map of the (structurally) sparse universe and tractable model checking.
@@ 2.1 Treewidth and the strong r-coloring numbers
Treewidth, as a measure of the global connectivity of a graph, plays an important role in graph theory. It was originally introduced by Bertelè and Brioschi under the name of dimension [10]. It was rediscovered by Halin [80], and again rediscovered by Robertson and Seymour [136] as part of their graph minors project. We refer to [13, 82] for surveys on various characterizations of treewidth and equivalent parameters and to [14, 29] for surveys on algorithmic applications.
@@
A tree decomposition of a graph $G$ is a pair $\mathcal{T}=(T,\mathrm{Bag})$, where $T$ is a tree and $\mathrm{Bag}\colon V(T)\to\mathcal{P}(V(G))$ is a mapping that assigns each node $x$ of $T$ to its bag $\mathrm{Bag}(x)\subseteq V(G)$ so that the following conditions are satisfied (see Figure 4):
@@
In the following, it will be convenient to assume that $T$ is a rooted tree. A rooted tree is a tree $T$ with a designated root vertex $r$. Assigning a root imposes the standard ancestor/descendant relation $\leq_{T}$ in $T$ where the root is the $\leq_{T}$ minimal element: a node $v$ is a descendant of all the nodes that appear on the unique path leading from $v$ to the root. Note that we treat a vertex as an ancestor and descendant of itself. The ancestors and descendants of $v$ excluding $v$ are its true ancestors and true descendants, respectively. The parent of a non-root node is its closest true ancestor. A rooted forest $F$ is a disjoint union of rooted trees. We write $\leq_{F}$ for the partial order that coincides with $\leq_{T}$ on each rooted tree $T$ in $F$.
@@
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a tree decomposition of $G$. This decomposition induces a quasi-order on $V(G)$ as follows. For $v\in V(G)$, let $x(u)$ be the unique $\leq_{T}$-minimal node of $T$ with $v\in\mathrm{Bag}(x)$. Define the quasi-order $\leq_{\mathcal{T}}$ on $V(G)$ by $u\leq_{\mathcal{T}}v$ if and only if $x(u)\leq_{T}x(v)$. Any linearization of $\leq_{\mathcal{T}}$ has very nice properties, which we will exploit later. In fact, we can associate a width to any linear order of $V(G)$ and thereby provide another characterization of treewidth.
@@
Let $G$ be a graph and let $\pi=(v_{1},\ldots,v_{n})$ be an order of $V(G)$. We construct the graph $G^{\pi}_{\textit{fill}}$, called the fill-in graph of $G$ with respect to the order $\pi$, by the following fill-in procedure. We let $G_{n}\coloneqq G$. Now, for $i=n-1,\ldots,1$, assume that the graph $G_{i+1}$ has been constructed. We define $G_{i}$ as the graph obtained from $G_{i+1}$ by adding all edges between non-adjacent neighbors of $v_{i+1}$ that are smaller than $v_{i}$. Finally, we let $G^{\pi}_{\textit{fill}}\coloneqq G_{1}$. It is immediate from the definition that $\pi$ is a perfect elimination order of $G^{\pi}_{\textit{fill}}$ (see Figure 5).
@@
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is strongly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}v$ for all internal vertices $w\in V(P)$ (see Figure 6). The set of vertices strongly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{SReach}_{r}[G,\pi,v]$. We define
@@ 2.2 Treedepth and the weak r-coloring numbers
Let $G$ be a graph. The treedepth $\operatorname{td}(G)$ of $G$ is the minimum depth of a rooted forest $F$ on the same vertex set as $G$ such that whenever $uv\in E(G)$, then $u\leq_{F}v$ or $v\leq_{F}u$ (see Figure 7).
@@
Let $G$ be a graph and $r$ a positive integer. Let $\pi$ be a linear order of $V(G)$. We say that a vertex $u\in V(G)$ is weakly $r$-reachable with respect to $\pi$ from a vertex $v\in V(G)$ if $u\leq_{\pi}v$ and there exists a path $P$ (possibly of length 00) between $u$ and $v$ of length at most $r$ with $w>_{\pi}u$ for all internal vertices $w\in V(P)$ (see Figure 8). The set of vertices weakly $r$-reachable by $v$ with respect to the order $\pi$ is denoted $\operatorname{WReach}_{r}[G,\pi,v]$. We define
@@ 2.4 Logic and transductions
A variable $x$ not in the scope of a quantifier is a free variable (we do not consider formulas with free second-order variables). A formula without free variables is a sentence. The quantifier rank $\mathrm{qr}(\varphi)$ of a formula $\varphi$ is the maximum nesting depth of quantifiers in $\varphi$. A formula without quantifiers is called quantifier-free.
@@
For every FO-formula $\varphi(\bar{x})$ and every class $\mathscr{C}$ of colored graphs with bounded treedepth there exists a quantifier-free formula $\tilde{\varphi}(\bar{x})$ and a linear time computable map $Y$ such that, for every $G\in\mathscr{C}$, $Y(G)$ is a guided expansion of $G$ such that for all tuples of vertices $\bar{v}$
@@ 3 General bounds for the generalized coloring numbers
In this section, we present lower and upper bounds for the generalized coloring numbers. It turns out that the numbers are intimately linked to the density of bounded-depth minors and bounded-depth topological minors. These are the central notions in the theory of bounded expansion and nowhere dense graph classes, which were introduced by Nešetřil and Ossona de Mendez [112, 114]. After introducing these concepts, we present simple lower bounds for the generalized coloring numbers in terms of the density of depth-$r$ minors, and depth-$r$ topological minors. Throughout this work when $r$ is not explictly quantified it denotes a non-negative integer.
@@ 3.1 Graph expansion and simple lower bounds
The following notions of bounded-depth minors and bounded-depth topological minors are the fundamental definitions in the theory of bounded expansion and nowhere dense graph classes introduced by Nešetřil and Ossona de Mendez [112, 114].
@@
Hence, expansion and topological expansion are functionally equivalent. The key observation for the proof is that in the branch set of a high-degree vertex in a bounded-depth minor model, we will always find a vertex of large degree. We can use such a vertex as the branching vertex of a high-degree vertex in a topological minor model.
@@
Bounded-depth minors in classes of bounded expansion have constant edge density by definition. It turns out that also bounded-depth minors in nowhere dense classes have small edge density.
@@ 3.2 Bounding r-admissibility
In view of Lemma 2.21, the algorithm of Theorem 3.15 also gives linear time computable approximations for the strong $r$-coloring numbers and weak $r$-coloring numbers. It was shown by Grohe et al. [77] that for $r\geq 3$, computing exact values for $\operatorname{wcol}_{r}(G)$ is NP-complete. Similarly, it is shown by Dvořák [49] that deciding whether a graph contains a depth-$1$ minor of density at least $d$ is NP-complete for $d\geq 2$ and by Muzi et al. [108] that deciding whether a graph contains a $1$-subdivision of density at least $d$ for $d\geq 2$ is also NP-complete. It was recently proved by Breen-McKay et al. [22] that determining if a graph has weak or strong $r$-coloring number at most $k$ is para-NP-hard when parameterized by $k$ for all $r\geq 2$. This answers an earlier question by the author and under standard complexity assumptions implies that there does not exist an algorithm that, given an $n$-vertex graph $G$, computes $\operatorname{col}_{r}(G)$ and $\operatorname{wcol}_{r}(G)$ in time $n^{f(r,\operatorname{col}_{r}(G))}$, or $n^{f(r,\operatorname{wcol}_{r}(G))}$, for any computable function $f$. Moreover, they showed that there exists a constant $c$ such that it is NP-hard to approximate the generalized coloring numbers within a factor of $c$.
@@ 3.4 Transitive fraternal augmentations
If $w$ is an $r$-fraternity function for every value of $r$, then $w$ is called a fraternity function (see Figure 12).
@@ 4 Tree-decomposable graphs
Let $\mathcal{T}=(T,\mathrm{Bag})$ be a rooted tree decomposition of $G$. For a non-root node $x$ with parent $y$, we define the adhesion of $x$ as $\mathrm{Adh}(x)=\mathrm{Bag}(x)\cap\mathrm{Bag}(y)$. If $x$ is the root, then we set $\mathrm{Adh}(x)=\emptyset$ by convention. The adhesion of the tree decomposition $\mathcal{T}=(T,\mathrm{Bag})$ is the maximum size of an adhesion set in $\mathcal{T}$, i.e., $\max_{x\in V(T)}|\mathrm{Adh}(x)|$. The torso of a node $x$ is the graph $\mathrm{Torso}(x)$ on vertex set $\mathrm{Bag}(x)$ where two vertices $u,v\in\mathrm{Bag}(x)$ are adjacent if and only if $uv\in E(G)$ or if there exists $y\neq x$ such that $u,v\in\mathrm{Bag}(y)$. Equivalently, $\mathrm{Torso}(x)$ is obtained from $G[\mathrm{Bag}(x)]$ by turning the adhesions of $x$ and of all children of $x$ into cliques. The margin of a node $x$ is the set $\mathrm{Margin}(x)=\mathrm{Bag}(x)\setminus\mathrm{Adh}(x)$. We say that $\mathcal{T}$ is a tree decomposition over a class $\mathscr{D}$ of graphs if $\mathrm{Torso}(x)\in\mathscr{D}$ for every node $x$ of $T$ (see Figure 13).
@@ 4.1 H-minor-free graphs
Let $G$ be a graph. A partition $\mathcal{P}=(V_{1},\ldots,V_{\ell})$ of $V(G)$ is a connected partition if each $V_{i}$ induces a connected subgraph of $G$. We denote the minor of $G$ obtained from $G$ by contracting each $V_{i}$ to a single vertex $v_{i}$ by $G/\mathcal{P}$.
Note that we care about the order of the $V_{i}$, that is, we consider $G/\mathcal{P}$ as an ordered graph. We inherit the above definitions of simplicial veritces, perfect elimination orders, etc. for connected partitions by identifying a part $V_{i}$ with the vertex $v_{i}$ in $G/\mathcal{P}$. For example, we call $V_{i}$ simplicial if $v_{i}$ is simplicial in $G/\mathcal{P}$ and the width of $\mathcal{P}$ is the width of the order $(v_{1},\ldots,v_{\ell})$ of $G/\mathcal{P}$.